Highlands Company uses the weighted-average method in its process costing system. It processes wood pulp for various manufacturers of paper products. Data relating to tons of pulp processed during June are provided below:  

Percent Completed
Tons of Pulp Materials Labor and Overhead
Work in process, June 1 80,600 84 % 25 %
Work in process, June 30 50,600 41 % 20 %
Started into production during June 300,600

Required:  

1. Compute the number of tons of pulp completed and transferred out during June.

2. Compute the equivalent units of production for materials and for labor and overhead for June.

  • Required 1
  • Required 2

Compute the number of tons of pulp completed and transferred out during June.

Tons completed and transferred out

Compute the equivalent units of production for materials and for labor and overhead for June.

Materials Labor and Overhead
Equivalent units of production
Answer #1

1.

Completed and transferred out = 80,600 + 300,600 - 50,600 = 330,600

2.

Equivalent units = Completed and transferred out + (% of completion of ending work in process)

Material = 330,600 + (50,600 X 41%) = 351,346

Labor and overhead = 330,600 + (50,600 X 20%) = 340,720